OpenStudy (anonymous):

You are pushing a 150-kg wooden crate in a straight line a distance of 4.5 m across a wooden floor at constant speed. The static and kinetic coefficients of friction are 0.42 and 0.30, respectively. What is the work done by friction on the crate?

OpenStudy (anonymous):

Use the kinetic coefficient of friction, as it's in motion. Work = Force*Distance, or \[\int\limits_{0}^{4.5} F dx\]

OpenStudy (p0sitr0n):

force = friction force. because it is moving, we use kinetic friction coefficient. thus, F = 0.3*m*g = 0.3*150kg*9.8m/s^2. W = f*x, W = F * 4.5m. thats it do the calculations for F plz, i dont have a calculator right now